OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

jimthompson5910 (jim_thompson5910):

what is the derivative of s(t)? what did you get?

OpenStudy (anonymous):

-2 - 6(x+h) ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you made an error somewhere

jimthompson5910 (jim_thompson5910):

that -2 should go away

OpenStudy (anonymous):

-2 - 6(x+h) -(-2-6x)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

then that over h?

jimthompson5910 (jim_thompson5910):

yes that's correct

jimthompson5910 (jim_thompson5910):

you'll find that the first -2 and the -(-2) will cancel

OpenStudy (anonymous):

is the answer -6?

OpenStudy (anonymous):

bc you get -2-6x-6h+2+6x/h = -6h/h = -6

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

s(t) = -2 - 6t s ' (t) = -6

jimthompson5910 (jim_thompson5910):

so no matter what t is, the instantaneous velocity will be -6

OpenStudy (anonymous):

ok thank you!

jimthompson5910 (jim_thompson5910):

you're welcome